# SLED: A Speculative LLM Decoding Framework for Efficient Edge Serving

### v2

@@ IV-C Cost Efficiency and System Throughput
The cost efficiency of token generation in edge computing scenarios is a critical factor and has been considered in various edge inference system designs [29, 30]. In this paper, we compare the proposed SLED system, a centralized serving system, and an all-edge decoding system in terms of both cost efficiency and whole-system token generation rate (WSTGR). To systematically analyze the trade-off between cost and performance, we construct Pareto front visualizations, which highlight the non-dominated configurations that achieve the best balance between monetary cost and system throughput.
In our experiments, the cost and throughput metrics for edge inference scenarios were rigorously computed based on a comprehensive capital expenditure (CAPEX) and operational expenditure (OPEX) model[31]. Specifically, we adopt the widely recognized CPU-hour cost model described by Walker[31] and the edge-compute modeling approach proposed by Eriksson[32]. The CAPEX component was determined by amortizing the purchase price of each edge device (Raspberry Pi 5 and Pi 4B, priced at \$80 [33] and \$75 [34] respectively) over a three-year lifetime, assuming an average device utilization rate of 70%. The OPEX component included electrical consumption calculated from experimentally measured average power draw (8 W for Pi 5 and 5 W for Pi 4B) [35] and industrial electricity rates (0.083 \$/kWh) [36]. Combining these costs, we obtained a unified hourly expense for each device, subsequently normalized by the experimentally measured token generation rates (tokens per second), as shown in Eq. 2. The resulting metric, expressed clearly as dollars per one thousand generated tokens (\$/1K tokens), enabled direct and transparent comparison across different experimental configurations and devices.
@@
Figure 6 compares following three deployment strategies along a common cost–performance plane. Specifically, the strategies are: 1) all-Server executes every token‐generation step on a bank of four NVIDIA A100–80 GB GPUs. 2) All-Edge places the same LLaMA draft model on each Raspberry Pi 5, with no server involvement. 3) SLED lets the Raspberry Pi 5 generates draft tokens, which are batch-verified on the A100 cluster of same configuration of centralized scenario. For every strategy we sweep two orthogonal factors: quantization precision (16-, 8-, and 4-bit) and edge-device count $N\in\{1,2,4,8,16\}$. Cost is monetised as dollars per one-thousand verified tokens, while throughput is already-defined WSTGR.
@@
We observe that SLED’s skyline consistently dominates the Pareto frontier, achieving lower cost per 1K verified tokens while sustaining higher overall throughput. For instance, at a budget of approximately \$0.13 per 1K tokens, SLED achieves a throughput of 137 tokens/s—representing a 3.5× improvement over the centralized baseline—while simultaneously reducing cost to just 29% of that incurred by the server-only configuration under comparable conditions. This advantage becomes more pronounced as the number of edge devices increases: whereas the All-Edge throughput saturates due to power constraints (e.g., increasing $N$ from 8 to 16 yields only marginal gains), SLED leverages server-side batching to amortize GPU usage, enabling near-linear throughput scaling beyond $N\geq 4$. Furthermore, quantization universally improves cost efficiency across all schemes by simultaneously reducing energy demand and increasing per-device generation rate. Notably, the 4-bit SLED configuration with 16 devices achieves 138 tokens/s at \$0.13 / 1K tokens, representing a 65% improvement in throughput over the best-performing All-Edge setup, with acceptable additional cost. These results substantiate the claim that SLED enables a superior cost–throughput trade-off, combining local cost efficiency with global throughput.
@@
Overall, the experimental evaluations underscore the significant advantages of SLED in distributed inference scenarios, including throughput, capacity, cost efficiency, showcasing insightful findings in the SLED system to motivate more explorations in the future work.
V Conclusion
@@
This paper presented the SLED framework, a novel distributed decoding solution designed for LLM deployment at the edge. Our extensive evaluation demonstrated that SLED significantly improves system throughput, capacity and cost efficiency compared to traditional centralized approaches. The integration of speculative local drafting and centralized verification establishes a balance of computational workload, making SLED particularly suitable for bringing LLMs towards the edge of network. It’s proved in this position paper that the SLED is more than a decoding enhancement—it opens the door to more foundational and elastic approach to resource-aware LLM serving at the edge.
VI Future Work
@@
Future research directions include exploring the smart queue and batching strategy on the edge server for latency-sensitive task and better server utilization. Additionally, enhancing the adaptive capabilities of SLED for dynamic environments, such as extending SLED’s applicability to multi-modal scenarios will be another interesting topic to focus on. Lastly, network condition and resource aware verification strategy could further broaden its practical impact in complex edge computing landscapes.

### v3

@@ IV-C Cost Efficiency and System Throughput
The cost efficiency of token generation in edge computing scenarios is a critical factor and has been considered in various edge inference system designs [30, 31]. In this paper, we compare the proposed SLED system, a centralized serving system, and an all-edge decoding system in terms of both cost efficiency and WSTGR. To systematically analyze the trade-off between cost and performance, we construct Pareto front visualizations, which highlight the non-dominated configurations that achieve the best balance between monetary cost and system throughput.
In our experiments, the cost and throughput metrics for edge inference scenarios were carefully computed based on a comprehensive capital expenditure (CAPEX) and operational expenditure (OPEX) model[32]. Specifically, we adopt the widely recognized CPU-hour cost model described by Walker[32] and the edge-compute modeling approach proposed by Eriksson[33]. The CAPEX component was determined by amortizing the purchase price of each edge device (Raspberry Pi 5 priced at \$80 [34]) over a three-year lifetime, assuming an average device utilization rate of 70%. The OPEX component included electrical consumption calculated from experimentally measured average power draw (8 W for Pi 5) [35] and industrial electricity rates (0.083 \$/kWh) [36]. Combining these costs, we obtained a unified hourly expense for each device, subsequently normalized by the experimentally measured token generation rates (tokens per second), as shown in Eq. 2. The resulting metric, expressed clearly as dollars per one thousand generated tokens (\$/1K tokens), enabled direct and transparent comparison across different experimental configurations and devices.
@@
Figure 6 compares the following three deployment strategies along a common cost–performance plane. Specifically, the strategies are: 1) all-Server executes every token‐generation step on a bank of four NVIDIA A100–80 GB GPUs. 2) All-Edge places the same LLaMA draft model on each Raspberry Pi 5, with no server involvement. 3) SLED lets the Raspberry Pi 5 generate draft tokens, which are batch-verified on the A100 cluster with the same configuration of the centralized scenario. For every strategy ,we sweep two orthogonal factors: quantization precision (16-, 8-, and 4-bit) and edge-device count $N\in\{1,2,4,8,16\}$. Cost is monetised as dollars per one-thousand verified tokens.
@@
We observe that SLED’s skyline consistently dominates the Pareto frontier, achieving lower cost per 1K verified tokens while sustaining higher overall throughput. For instance, with the same system capacity and quantization level, SLED achieves a throughput of 137 tokens/s—representing a 3.5× improvement over the centralized baseline—while simultaneously reducing cost to just 29% of that. This advantage becomes more pronounced as the number of edge devices increases. Furthermore, quantization universally improves cost efficiency across all schemes by simultaneously reducing energy demand and increasing per-device generation rate. Notably, the 4-bit SLED configuration with 16 devices achieves 137 tokens/s at \$0.13 / 1K tokens, representing a 65% improvement in throughput over the best-performing All-Edge setup, with acceptable additional cost. These results substantiate the claim that SLED enables a superior cost–throughput trade-off, combining local cost efficiency with global throughput.
@@
Overall, the experimental evaluations underscore the significant advantages of SLED in distributed inference scenarios, including throughput, capacity, and cost efficiency, showcasing insightful findings in the SLED system to motivate more explorations in future work.
V Conclusion and Future Work
@@
This position paper presented the SLED, a novel distributed decoding framework designed for LLM deployment at the edge. Our extensive evaluation demonstrated that SLED significantly improves system throughput, capacity, and cost efficiency compared to traditional centralized approaches. The integration of speculative local drafting and centralized verification establishes a balance of computational workload, making SLED particularly suitable for bringing LLMs towards the edge of the network. It’s highlighted in this position paper that the SLED is more than a decoding enhancement—it opens the door to a more foundational and elastic approach to resource-aware LLM serving at the edge.
@@
Future research directions include exploring the adaptive queue and batching strategy on the edge server for latency-sensitive tasks and better server utilization. Additionally, enhancing the adaptive capabilities of SLED for dynamic environments, such as extending SLED’s applicability to multi-modal scenarios, will be another interesting topic to focus on. Lastly, network conditions and resource-aware verification strategy could further broaden its practical impact in complex edge computing landscapes.